Anne and Nancy use a metal alloy that is 15% copper to make jewelry. How many ounces of 10% alloy must be mixed with a 21% alloy to form 121 ounces of the desired alloy?
Answer by robertb(5830)

x = #oz. of 10% alloy,
121-x = #oz. of 21% alloy. Then
0.10x + 0.21(121 - x) = 0.15*121,
0.10x + 25.41-0.21x = 18.15,
-0.11x = -7.26,
x = 66 oz, amount of 10% alloy,
121 - x = 55 oz, amount of 21% alloy.